OpenStudy (anonymous):

Find using quotient rule: j'(x) = [(-1/4)x+1]/[-2(abs(x-2))+4] (simpler looking equation in comments)

OpenStudy (anonymous):

\[j'(X)=\frac{ \frac{ -1 }{ 4 }x+1 }{ -2\left| x-2 \right|+4 }\]

OpenStudy (anonymous):

I know the formula for the quotient rule, it's just the absolute value that's throwing me off

OpenStudy (anonymous):

anyone? :/

OpenStudy (anonymous):

so would the derivative of the entire denominator simplify to -2x?

OpenStudy (anonymous):

ok! i think i got it. It says i got the problem right now at least. thanks for the help!

OpenStudy (ranga):

Hold on. I may have to flip them. Derivative of |x| is x / |x| Derivative of |x - 2| = (x - 2) / |x - 2|

OpenStudy (anonymous):

Got it. Thanks a lot that really helped!

OpenStudy (ranga):

This link may help too: http://www.youtube.com/watch?v=-z0Kii-yU1Q